What amount of each mixture, one 95% alcohol and the other 15% alcohol, must be used to make 10 liters of a mixture which is 45% alcohol.
How would you express this problem in an equation?

Found 2 solutions by scott8148, stanbon:
Answer by scott8148(6628)

let x="liters of 95%", so 10-x="liters of 15%"

(x)(95%) + (10-x)(15%) = (10)(45%) ___ multiply by 100 to clear percentages

95x + 150 - 15x = 450

80x = 300

x = 3.75

10-x = 6.25

Answer by stanbon(57387)

What amount of each mixture, one 95% alcohol and the other 15% alcohol, must be used to make 10 liters of a mixture which is 45% alcohol.
--------------------------
Equation:
alcohol + alcohol = alcohol
0.95x + 0.15(10-x) = 0.45*10
----
Multiply thru by 100 to get:
95x + 15*10 - 15x = 45*10
80x = 30*10
x = (1/8)30
x = 3.75 liters (amt. of 95% solution in the mixture)
10-3.75 = 6.25 liters (amt. of 15% solution in the mixture)
==================================
Cheers,
Stan H.